OpenStudy (anonymous):

how to expand [1-2x]^11

OpenStudy (apoorvk):

\[(a+b)^n = \sum_{r=0}^{n} * ^{n}C _{r}^{} a^rb ^{n-r}\] just remove that "*".

OpenStudy (apoorvk):

so here, a=1 b= (-2x)

OpenStudy (anonymous):

i dont know

OpenStudy (cwrw238):

this is a binomial expansion (1 - 2x)^11 = 1 + 11C1 (-2x) + 11C2 (-2x)^2 + 11C3 (-2x)^3+ ..........+ (-2x)^11

OpenStudy (cwrw238):

do u know what 11C3 means and how to calculate it

OpenStudy (cwrw238):

11C3 means number of combinations of 3 from 11, 11C4 means number of 4 from 11 etc